x+y+3z=-4
-x-y-2z=5
2x-z=-3
Work being shown

Answer by Alan3354(69443)

x+y+3z=-4
-x-y-2z=5
2x-z=-3
============
Add Eqn 1 and Eqn 2.
etc.